# FaceCoT: A Benchmark Dataset for Face Anti-Spoofing with Chain-of-Thought Reasoning

### v1

@@ 4 Methodology
In FAS tasks, the ability to discern fine-grained facial details is critical to model performance. The FaceCoT dataset provides rich fine-grained visual descriptions of facial features, which can effectively guide the learning of a powerful visual encoder. However, adopting a single-stage training strategy typically results in a suboptimal visual encoder. This is because the model is required to learn both reasoning and classification tasks simultaneously, and the faster convergence of the classification objective often leads to insufficient optimization of the reasoning component. Consequently, the model cannot fully leverage the fine-grained visual cues in the CoT annotations, preventing the visual encoder from capturing critical spoofing artifacts and thereby capping overall performance. To fully leverage CoT data, we propose a CoT-Enhanced Progressive Learning (CEPL) strategy. As shown in Figure 4, CEPL consists of two stages: (1) Visual Enhancement Pre-training, we fine-tune exclusively on CoT annotations—using language-guided supervision to sharpen the vision encoder’s sensitivity to subtle facial features; (2) Multi-task Joint Training, we jointly optimize CoT reasoning and spoof classification to achieve synergy between visual understanding and logical inference.
@@ 4.2 Multi-task Joint Training
The objective of this stage is to achieve synergistic enhancement of reasoning and classification capabilities. To this end, we employ a joint training strategy over both CoT reasoning and binary classification tasks. During model initialization, we retain the vision encoder from Visual Enhancement Pre-training to preserve its fine-grained facial representations, while restoring all other modules to their original pretrained weights. We then insert lightweight LoRA hu2022lora adapters into the LLM for targeted fine-tuning. Training proceeds on a combined dataset of CoT-annotated samples and binary labels, guiding the network to balance cross-modal reasoning with accurate spoofing detection. After this stage, the model not only reliably distinguishes real from spoofing faces but also produces coherent CoT explanations for its decisions.
Table 1: Comparison of evaluation metrics against state-of-the-art methods on 11 benchmark FAS datasets. All baselines ViTAF ; CLIP ; FLIP ; I-FAS are trained on CelebA-Spoof; “Ours-CelebA” is trained on the CelebA-Spoof dataset CelebA-Spoof annotated by the FAS caption model, “Ours-100K” on FaceCoT-Gold100K, and “Ours-All” on the full FaceCoT dataset.
@@ 5.1 Experimental settings
To validate the robustness and generalization of our data, we use FaceCoT as the source domain and perform cross-domain testing on 11 other datasets as target domains. These datasets include MSU-MFSD msu-mfsd , CASIA-MFSD casia-mfsd , Idiap Replay Attack replay-attack , OULU-NPU oulu-npu , SIW siw , Rose-Youtu rose-youtu , HKBU-MARs-V1+ hkbu , WMCA WMCA , SIW-M-V2 siw-mv2 , CASIA-SURF-3DMask surf-3d-mask , and HiFiMask hifi-mask . For a fair comparison with previous methods I-FAS , we also replicate their training setup by using only the annotated CelebA-Spoof CelebA-Spoof subset as the source domain while testing on the same 11 benchmarks.
@@
We resize all images to $448\times 448\times 3448 × 448 × 3$ with RGB channels. We choose MiniCPMV-2.6-8B minicpm for its lightweight multimodal architecture, which exhibits strong cross-modal fusion capabilities across diverse scenarios. These characteristics make it particularly well-suited to the core demands of FAS tasks, where effective integration of multimodal features is essential. We employ the AdamW optimizer with an initial learning rate of 1e-6 and a weight decay of 0.1. Training is conducted for up to 10 epochs with a global batch size of 256, distributed across eight NVIDIA A100 GPUs. As in previous works I-FAS ; FLIP ; ViTAF , we use Half Total Error Rate (HTER) and Area Under the Receiver Operating Characteristic Curve (AUC) as evaluation metrics.
@@ 5.2 Results analysis
Analyzing the results in Table 1, we can draw three conclusions: (1) Overall performance gain: Training on FaceCoT-All yields an average AUC increase of 4.06% and an HTER reduction of 5% over the previous SOTA methods I-FAS ; ViTAF ; CLIP ; FLIP , and it achieves the best score on nearly every evaluation set. (2) Cross-domain generalization: HKBU-MARs-V1+ hkbu and HiFiMask hifi-mask include spoofing types absent from the source data—e.g., transparent, plaster, and resin masks. Despite this severe distribution shift, our approach surpasses the previous best by roughly 10% on HKBU-MARs-V1+ and 14% on HiFiMask, demonstrating strong robustness to unseen attack modalities. (3) Single-source comparison: When training is restricted to CelebA-Spoof CelebA-Spoof alone, the model still outperforms the previous SOTA: average HTER drops a further 2.74% and AUC rises 2.66%, again securing the top scores on HKBU-MARs-V1+ and Rose-Youtu rose-youtu . This confirms that the CoT annotations and the CEPL training strategy remain effective even under limited-source settings.
@@
Methods Stage-1 Stage-2 Stage-3 Results
@@ 6 Ablation study
In this section, we systematically evaluate each major contribution of our framework. All ablation experiments are trained on the FaceCoT-Gold100K subset to ensure a consistent comparison baseline.
@@
To validate the soundness of our proposed CEPL method, we perform two ablation studies: (1) To validate the effectiveness of the proposed CEPL method, we compare it with the single-stage Multi-task Joint Training method. (2) Meanwhile, we further extend the CEPL framework by applying RL VRFT after the second stage, in order to investigate whether RL can bring additional performance gains. As shown in Table 2, we make two key observations: (1) We find that the CEPL method outperforms single-stage training with an increase of 0.68% in AUC and a decrease of 1.19% in HTER. This indicates that stage-wise optimization helps the model better capture the knowledge transfer relationship between different tasks, effectively reducing task interference and improving overall performance. (2) The performance metrics after applying RL are similar to the original method, but do not surpass it. We hypothesize that in scenarios with limited SFT data, initializing the model with a small amount of SFT and applying RL to unlabeled data can significantly improve performance compared to relying solely on SFT. However, in our study, since the model has already undergone extensive SFT using a large amount of high-quality data, the marginal gains from applying RL afterward are notably reduced.
Table 3: Ablation study on CoT data, with comparison to a setup using only binary label data.
@@
Table 4: Ablation study on resolution, comparing input sizes of $\times×224$ and $\times×448$.
@@
$224\times 224224 × 224$ 11.28 94.05
$448\times 448448 × 448$ 8.84 95.91
Figure 5: The figure shows the outputs of different FAS methods: (a) Traditional binary classification method; (b) Other MLLMs I-FAS can answer classifications and provide simple descriptions; (c) Our method can not only answer classification questions, but also provide systematic reasoning analysis.
@@
To validate the effectiveness of the CoT data that we construct during training, we design an ablation experiment at a resolution of $448\times 448448 × 448$, comparing single-stage training using only binary-labeled data against training with CoT data under the CEPL framework. The experimental results are shown in Table 4. It can be seen that the model trained with CoT data achieves an HTER of 7.65% and an AUC of 96.59%, both outperforming the model trained without CoT annotations. This result indicates that CoT data provides the model with richer intermediate process information, helping the model better understand complex tasks and significantly enhancing its generalization ability and robustness. Moreover, Figure 5 illustrates that the CoT-trained model not only makes more accurate classification but also exposes its full CoT—especially on attack samples—providing interpretable rationales that further strengthen its cross-domain reliability. This combined evaluation confirms both the performance and interpretability benefits of our constructed CoT dataset.
@@
Capturing fine-grained visual features is critical for FAS detection, and higher input resolution typically provides richer local texture information, thereby enhancing the model’s visual representation capacity. To systematically assess the impact of input resolution on model performance, we design a comparative experiment with different resolutions, keeping all other settings consistent. Specifically, we set the input resolutions to $224\times 224224 × 224$ and $448\times 448448 × 448$, respectively, and compare the performance of the single-stage trained model at both resolutions. The experimental results are shown in Table 4. From the results, it can be seen that when the resolution is $224\times 224224 × 224$, the model’s performance on all metrics is slightly lower than at the $448\times 448448 × 448$ resolution. This suggests that a higher resolution helps the model capture richer detailed features, further enhancing performance.
@@ 7 Conclusions
We introduce FaceCoT, the first large-scale VQA dataset for FAS with 1.08 million detailed CoT annotations that cover a wide spectrum of attack types. The dataset begins with 100K high-quality samples, then is expanded to 1.08 million using a reinforcement learning-enhanced caption model. Additionally, we propose a CEPL method that achieves effective synergy between semantic guidance and attack type discrimination. Extensive experiments confirm both the value of the FaceCoT and the effectiveness of our method, delivering an average improvement of 4.06% over current state-of-the-art approaches. We believe that the open-source release of the FaceCoT dataset not only provides a valuable resource for the research community but also lays a solid data foundation and methodological reference for building stronger and more trustworthy FAS systems.
Appendix A FaceCoT dataset: additional details
A.1 GPT-4o annotation details
@@
[htbp] System Prompt In a face anti-spoofing detection system, a person has submitted an image for inspection. Now please assume the role of the examiner and respond to the question. I need you to strictly follow the format containing six specific parts: caption, facical description, facial attributes, reasoning, spoofing description, and conclusion. It is crucial that you adhere exactly to the outlined structure, and that the final answer in conclusion matches the standard correct answer exactly. Further explanation: In caption, provide a holistic description of the image. In facial description, focus on describing the facial features of the person in the image. In facial attributes, list the facial attributes. In reasoning, perform step-by-step logical reasoning based on the image description, facial description, and facial attributes. In spoofing description, describe the spoofing method and highlight the critical spoofing characteristics in detail. In conclusion, give the final answer “Yes” or “No” that must match the standard answer exactly. Format as follows: <Caption> [Holistic description of the image] </Caption> <Facial Description> [Detailed description of the person’s facial features] </Facial Description> <Facial Attributes> [Describe facial attributes: "eyes": ..., "mouth": ...etc.] </Facial Attributes> <Reasoning> [Step-by-step logical reasoning based on the above] </Reasoning> <Spoofing Description> [Details of the spoofing method: ...] </Spoofing Description> <Conclusion> [Answer “Yes” or “No” matching the standard answer exactly] </Conclusion> Please carefully apply this format to analyze the given image and answer the related question.
Furthermore, to prevent the model from detecting spoofing-related features without being able to establish clear decision boundaries (example of such ambiguous case is illustrated in Figure 7), we provide tailored hints for samples of different spoofing types to guide the model toward correct reasoning and conclusions, as presented in Table 5. For spoofing types that the model can readily distinguish, no hints are provided.
Figure 7: Example of a decision-boundary failure: the model correctly detects newspaper features but does not classify the image as spoofing (top). After adding a hint (bottom), the model refines its reasoning and accurately identifies the image as a spoof, demonstrating improved CoT guidance with better decision boundaries.
@@
A.2 Hard case handling
In the main text, we describe that, following GPT-4o annotation, we apply a regular expression to extract the result within the "<Conclusion>...</Conclusion>" tag and cross-check it against the original label; any mismatch is deemed an annotation failure. Samples that could not be correctly labeled after two annotation rounds are designated as “hard cases,” resulting in a total of 581 instances. These hard cases are then corrected by human experts. Figure 8 illustrates one such example: experts first verify whether the conclusion is correct, then diagnose why the reasoning and spoofing description are inaccurate; if the reasoning is flawed, they replace it with a correct step-by-step rationale; if the visual feature description is inconsistent, they refine it and iteratively update the subsequent reasoning. Through this expert review and correction process, we ultimately obtain the complete, high-quality FaceCoT-Gold100K dataset. Similarly, samples in the data expansion that fail to be correctly annotated by our caption model are also reviewed and corrected by human experts, leading to the construction of the complete FaceCoT-Silver982K dataset.
@@ A.2 Hard case handling
A.3 Statistics
Our FaceCoT dataset comprises two subsets—FaceCoT-Gold100K and FaceCoT-Silver982K—and encompasses living faces alongside 14 distinct spoofing attack types. Here, we present the relative proportions of types across the two datasets in Figure 9 and report the exact sample counts for every category in both subsets in Table 6. Examples of each attack category are illustrated in Figure 10
Table 6: Number of samples per category in the FaceCoT-Gold100K and FaceCoT-Silver982K subsets
@@ A.3 Statistics
Appendix B Methodology details
B.1 Reinforcement Learning in training caption model
@@
We design a dual-reward scheme targeting both semantic accuracy and output format compliance:
@@
  • Format compliance reward: We verify whether the model’s output strictly follows the prescribed FaceCoT template. If the structural format is correct, the reward is 1; otherwise, 0.
This dual-reward scheme simultaneously enforces correct annotation content and adherence to the FaceCoT formatting guidelines.
@@
We initialize the policy model with a version pre-trained via Supervised Fine-Tuning (SFT). Given an input image and its associated task prompt, the policy model generates a CoT response. Each response is scored according to the dual-reward functions above, and the resulting reward signal is used to update the policy via RL. To stabilize training, we employ the SFT model as a fixed reference: we compute the KL divergence between the policy’s output distribution and that of the reference model, using it as a penalty term to prevent the policy from drifting too far from its initial semantic space. This balance preserves output reliability while enabling effective exploration.
@@
To enhance the caption model’s annotation capability and task adaptability on unseen data, we directly use the unlabeled images from the target annotation corpus as input during the RL stage. This construction endows the caption model with strong task-specific adaptation.
@@
We first randomly sample 2,000 instances from the unlabeled dataset to construct a test set for evaluation. Then, we use two models to perform automatic annotation on this test set: one trained solely with SFT, and the other further optimized with RL based on the SFT model. From the generated outputs, we extract the result within the "<Conclusion>...</Conclusion>" tags and compare them with the original labels of the samples. If the two labels match exactly, the annotation is considered correct; otherwise, it is considered incorrect. The final annotation accuracy is calculated using the following formula:
[…]
B.2 Reinforcement Learning in CoT-Enhanced Progressive Learning (CEPL) framework

### v2

@@ 4 Methodology
In FAS tasks, the ability to discern fine-grained facial details is critical to model performance. The FaceCoT dataset provides rich fine-grained visual descriptions of facial features, which can effectively guide the learning of a powerful visual encoder. However, adopting a single-stage training strategy typically results in a suboptimal visual encoder. This is because the model is required to learn both reasoning and classification tasks simultaneously, and the faster convergence of the classification objective often leads to insufficient optimization of the reasoning component. Consequently, the model cannot fully leverage the fine-grained visual cues in the CoT annotations, preventing the visual encoder from capturing critical spoofing artifacts and thereby capping overall performance. To fully leverage CoT data, we propose a CoT-Enhanced Progressive Learning (CEPL) strategy. As shown in  Fig. 4, CEPL consists of two stages: (1) Visual Enhancement Pre-training, we fine-tune exclusively on CoT annotations, which uses language-guided supervision to sharpen the vision encoder’s sensitivity to subtle facial features; (2) Multi-task Joint Training, we jointly optimize CoT reasoning and spoof classification to achieve a stronger synergy between visual understanding and logical inference.
@@ 4.2 Multi-task Joint Training
The objective of this stage is to achieve synergistic enhancement of reasoning and classification capabilities. To this end, we employ a joint training strategy over both CoT reasoning and binary classification tasks. During model initialization, we retain the vision encoder from Visual Enhancement Pre-training to preserve its fine-grained facial representations, while restoring all other modules to their original pretrained weights. We then apply LoRA [13] modules to the LLM for targeted fine-tuning. Training proceeds on a combined dataset of CoT-annotated samples and binary labels, guiding the network to balance cross-modal reasoning with accurate spoofing detection. After this stage, the model not only reliably distinguishes real from spoofing faces but also produces coherent CoT explanations for its decisions.
Table 1: Comparison of evaluation metrics against state-of-the-art methods on 11 benchmark FAS datasets. All baselines are trained on CelebA-Spoof; “Ours-CelebA” is trained on the CelebA-Spoof dataset annotated by the FAS caption model, “Ours-100K” on FaceCoT-Gold100K, and “Ours-All” on the full FaceCoT dataset.
@@ 5.1 Experimental settings
To validate the robustness and generalization of our data, we use FaceCoT as the source domain and perform cross-domain testing on 11 other datasets as target domains. These datasets include MSU-MFSD [38], CASIA-MFSD [46], Idiap Replay Attack [6], OULU-NPU [3], SIW [23], Rose-Youtu [16], HKBU-MARs-V1+ [22], WMCA [10], SIW-M-V2 [12], CASIA-SURF-3DMask [42], and HiFiMask [20]. For a fair comparison with previous methods [43], we also replicate their training setup by using only the annotated CelebA-Spoof [45] subset as the source domain while testing on the same 11 benchmarks.
@@
We resize all images to $448\times 448\times 3$ with RGB channels. We choose MiniCPMV-2.6-8B [40] as the backbone VLM for its lightweight multimodal architectureand strong cross-modal fusion capabilities across diverse scenarios; results with alternative backbone VLMs are provided in Appendix. These characteristics make it particularly well-suited to the core demands of FAS tasks, where effective integration of multimodal features is essential. We employ the AdamW optimizer with an initial learning rate of 1e-6 and a weight decay of 0.1. Training is conducted for up to 10 epochs with a global batch size of 256, distributed across 8 NVIDIA A100 GPUs. To ensure robustness, we run all experiments with three different random seeds and report the averaged results. As in previous works [43, 31, 14], we use Half Total Error Rate (HTER) and Area Under the Receiver Operating Characteristic Curve (AUC) as evaluation metrics.
@@ 5.2 Results analysis
Analyzing the results in  Tab. 1, we can draw three conclusions: (1) Overall performance gain: Training on FaceCoT-All yields an average AUC increase of 4.06% and an HTER reduction of 5.00% over the previous state-of-the-art methods, and it achieves the best score on every evaluation set. (2) Cross-domain generalization: HKBU-MARs-V1+ and HiFiMask include spoofing types absent from the source data(e.g., transparent, plaster, and resin masks). Despite this severe distribution shift, our approach surpasses the previous best AUC by roughly 10% on HKBU-MARs-V1+ and 14% on HiFiMask, demonstrating strong robustness to unseen attack modalities. (3) Single-source comparison: When training is restricted to CelebA-Spoof alone, the model still outperforms the previous state-of-the-art methods: average HTER drops a further 2.74% and AUC rises 2.66%, again securing the top scores on HKBU-MARs-V1+ and Rose-Youtu. This confirms that the CoT annotations and the CEPL training strategy remain effective even under limited-source settings.
@@
Methods Stage1 Stage2 Stage3 Results
@@ 6 Ablation study
In this section, we systematically evaluate each major contribution of our framework. All ablation studies are trained on the FaceCoT-Gold100K subset to ensure a consistent comparison baseline.
@@
To validate the soundness of our proposed CEPL method, we perform two ablation studies: (i) To validate the effectiveness of the proposed CEPL method, we compare it with the single-stage Multi-task Joint Training method. (ii) Meanwhile, we further extend the CEPL framework by applying RL [26] after the second stage, in order to investigate whether RL can bring additional performance gains. As shown in  Tab. 2, we make two key observations: (1) We find that the CEPL method outperforms single-stage training with an increase of 0.68% in AUC and a decrease of 1.19% in HTER. This indicates that stage-wise optimization helps the model better capture the knowledge transfer relationship between different tasks, effectively reducing task interference and improving overall performance. (2) The performance metrics after applying RL are similar to the original method. We hypothesize that in scenarios with limited SFT data, initializing the model with a small amount of SFT and applying RL to unlabeled data can significantly improve performance compared to relying solely on SFT. However, in our study, the model has already undergone extensive SFT with a large amount of high-quality data, resulting in strong baseline performance, so the marginal gains from applying RL afterward are notably reduced.
Figure 5: The figure shows the outputs of different FAS methods: (a) Traditional binary classification method; (b) Other MLLMs [43] can answer classifications and provide simple descriptions; (c) Our method can not only answer classification questions, but also provide systematic reasoning analysis.
Table 3: Ablation study on CoT data, with comparison to a setup using only binary label data of FaceCoT-Gold100K.
@@
Table 4: Ablation study on resolution, comparing input sizes of $\times224$ and $\times448$.
@@
$224\times 224$ 11.28 94.05
$448\times 448$ 8.84 95.91
@@
To validate the effectiveness of the CoT data that we construct during training, we design an ablation experiment at a resolution of $448\times 448$, comparing single-stage training using only binary-labeled data against training with CoT data under the CEPL framework. The experimental results are shown in  Tab. 3. It can be seen that the model trained with CoT data achieves an HTER of 7.65% and an AUC of 96.59%, both outperforming the model trained without CoT annotations. This result indicates that CoT data provides the model with richer intermediate process information, helping the model better understand complex tasks and significantly enhancing its generalization ability and robustness. Moreover, Figure 5 illustrates that the CoT‐trained model not only makes more accurate classification but also exposes its full CoT, especially on attack samples, providing interpretable rationales that further strengthen its cross-domain reliability. This combined evaluation confirms both the performance and interpretability benefits of our constructed CoT dataset.
@@
Capturing fine-grained visual features is critical for FAS detection, and higher input resolution typically provides richer local texture information, thereby enhancing the model’s visual representation capacity. To systematically assess the impact of input resolution on model performance, we design a comparative experiment with different resolutions, keeping all other settings consistent. Specifically, we set the input resolutions to $224\times 224$ and $448\times 448$, respectively, and compare the performance of the single-stage trained model at both resolutions. The experimental results are shown in  Tab. 4. From the results, it can be seen that when the resolution is $224\times 224$, the model’s performance on all metrics is slightly lower than at the $448\times 448$ resolution. This suggests that a higher resolution helps the model capture richer detailed features, further enhancing performance.
@@ 7 Conclusions
We introduce FaceCoT, the first large-scale VQA benchmark dataset for FAS with 1.08 million detailed CoT annotations that cover a wide spectrum of attack types. The dataset begins with 100K high-quality samples, then is expanded to 1.08 million using a reinforcement learning-enhanced caption model. Additionally, we propose a CEPL method that achieves effective synergy between semantic guidance and attack type discrimination. Extensive experiments confirm both the value of the FaceCoT and the effectiveness of our method, delivering an average AUC improvement of 4.06% over current state‑of‑the‑art approaches. We believe that the open-source release of the FaceCoT dataset not only provides a valuable resource for the research community but also lays a solid data foundation and methodological reference for building stronger and more trustworthy FAS systems.
Supplementary Material
1 FaceCoT dataset: additional details
1.1 GPT-4o annotation details
@@
System Prompt In a face anti-spoofing detection system, a person has submitted an image for inspection. Now please assume the role of the examiner and respond to the question. I need you to strictly follow the format containing six specific parts: caption, facical description, facial attributes, reasoning, spoofing description, and conclusion. It is crucial that you adhere exactly to the outlined structure, and that the final answer in conclusion matches the standard correct answer exactly. Further explanation: In caption, provide a holistic description of the image. In facial description, focus on describing the facial features of the person in the image. In facial attributes, list the facial attributes. In reasoning, perform step-by-step logical reasoning based on the image description, facial description, and facial attributes. In spoofing description, describe the spoofing method and highlight the critical spoofing characteristics in detail. In conclusion, give the final answer “Yes” or “No” that must match the standard answer exactly. Format as follows: <Caption>[Holistic description of the image] </Caption><Facial Description>[Detailed description of the person’s facial features] </Facial Description><Facial Attributes>[Describe facial attributes: ”eyes”: …, ”mouth”: …etc.] </Facial Attributes><Reasoning>[Step-by-step logical reasoning based on the above] </Reasoning><Spoofing Description>[Details of the spoofing method: …] </Spoofing Description><Conclusion>[Answer “Yes” or “No” matching the standard answer exactly] </Conclusion>Please carefully apply this format to analyze the given image and answer the related question.
Furthermore, to prevent the model from detecting spoofing‐related features without being able to establish clear decision boundaries (example of such ambiguous case is illustrated in Figure 7), we provide tailored hints for samples of different spoofing types to guide the model toward correct reasoning and conclusions, as presented in Table 5. For spoofing types that the model can readily distinguish, no hints are provided.
Figure 7: Example of a decision‐boundary failure: the model correctly detects newspaper features but does not classify the image as spoofing (top). After adding a hint (bottom), the model refines its reasoning and accurately identifies the image as a spoof, demonstrating improved CoT guidance with better decision boundaries.
@@
1.2 Hard case handling
In the main text, we describe that, following GPT-4o annotation, we apply a regular expression to extract the result within the "<Conclusion>...</Conclusion>" tag and cross-check it against the original label; if the match fails, this annotation is regarded as failed. Samples that could not be correctly labeled after two annotation rounds are designated as “hard cases,” resulting in a total of 581 instances. These hard cases are then corrected by human experts. Figure 8 illustrates one such example: experts first verify whether the conclusion is correct, then diagnose why the reasoning and spoofing description are inaccurate; if the reasoning is flawed, they replace it with a correct step-by-step rationale; if the visual feature description is inconsistent, they refine it and iteratively update the subsequent reasoning. Through this expert review and correction process, we ultimately obtain the complete, high-quality FaceCoT-Gold100K dataset. Similarly, samples in the data expansion that fail to be correctly annotated by our caption model are also reviewed and corrected by human experts, leading to the construction of the complete FaceCoT-Silver982K dataset.
@@ 1.2 Hard case handling
1.3 Statistics
Our FaceCoT dataset comprises two subsets—FaceCoT-Gold100K and FaceCoT-Silver982K—and encompasses living faces alongside 14 distinct spoofing attack types. Here, we present the relative proportions of types across the two datasets in Figure 9 and report the exact sample counts for every category in both subsets in Table 6. Examples of each attack category are illustrated in Figure 10. For the annotation of FaceCoT-Gold100K, we used the GPT-4o API, incurring an average cost of approximately \$0.01 per image. To enable large-scale annotation at lower cost, we further trained a caption model via SFT and RL, which required 8 NVIDIA A100 GPUs for about one day, and subsequently employed this model to annotate FaceCoT-Silver982K at a total cost of roughly 288 GPU-hours. Finally, in the human refinement stage, six annotators manually cleaned and verified samples from both subsets over a period of three days.
Table 6: Sample counts per category in the FaceCoT-Gold100K and FaceCoT-Silver982K subsets
@@ 1.3 Statistics
2 Methodology details
2.1 Reinforcement Learning in training caption model
@@
We design a dual‐reward scheme targeting both semantic accuracy and output format compliance:
@@
  • Format compliance reward: We verify whether the model’s output follows the prescribed FaceCoT template. If the structural format is correct, the reward is 1; otherwise, 0.
This dual‐reward scheme simultaneously enforces correct annotation content and adherence to the FaceCoT formatting guidelines.
@@
We initialize the policy model with a version pre‐trained via SFT. Given an input image and its associated task prompt, the policy model generates a CoT response. Each response is scored according to the dual-reward functions above, and the resulting reward signal is used to update the policy via RL. To stabilize training, we employ the SFT model as a fixed reference: we compute the KL divergence between the policy’s output distribution and that of the reference model, using it as a penalty term to prevent the policy from drifting too far from its initial semantic space. This balance preserves output reliability while enabling effective exploration.
@@
To enhance the caption model’s annotation capability and task adaptability on unseen data, we directly use the unlabeled images from the target annotation corpus as input during the RL stage. This construction endows the caption model with strong task‐specific adaptation.
@@
We first randomly sample 2,000 instances from the dataset that have not been annotated to construct a test set for evaluation. Then, we use two models to perform automatic annotation on this test set: one trained solely with SFT, and the other further optimized with RL based on the SFT model. From the generated outputs, we extract the result within the "<Conclusion>...</Conclusion>" tags and compare them with the original labels of the samples. If the two labels match exactly, the annotation is considered correct; otherwise, it is considered incorrect. The final annotation accuracy(ACC) is calculated using the following formula:
$\text{ACC}=\frac{\text{Count}(\texttt{conclusion}=\texttt{label})}{\text{Count}(\texttt{conclusion}=\texttt{label})+\text{Count}(\texttt{conclusion}\neq\texttt{label})}$ (1)
2.2 Reinforcement Learning in CoT-Enhanced Progressive Learning (CEPL)